# Token Spammers, Rug Pulls, and SniperBots: An Analysis of the Ecosystem of Tokens in Ethereum and the Binance Smart Chain (BNB)

### v2

@@ 4.1 The Token Dataset
To build the Token dataset, we perform multiple steps. We first collect all the contract creation transactions on both blockchains, then we expand our dataset by collecting all the contracts that emit a Transfer event. This process is necessary to find tokens generated by internal transactions. Finally, we refine our dataset by selecting only the smart contracts compliant with the ERC-20 (resp. BEP-20) standard.
@@ 4.1.1 Gathering Smart Contracts
As a first step to building the Token dataset, we collect all the contract creation transactions issued by EOAs. As mentioned in Section 2, EOAs can deploy a smart contract by sending a contract creation transaction to the zero address. We process all the transactions in the considered time frame in BNB Smart Chain (2.6 billion transactions) and Ethereum (1.4 billion transactions). We collect 2,195,399 and 4,420,389 contract creation transactions respectively.
However, tokens can also be created by a smart contract itself. Indeed, it could be the case that an EOA calls a smart contract method, and its execution generates a new ERC-20 (or BEP-20) compliant smart contract. In this case, the token is created with a so-called internal transaction. Despite the name, internal transactions are not real transactions but instead calls performed by smart contracts. These kinds of transactions are stored off-chain—they are not visible simply by parsing the blockchain.
To track tokens created by internal transactions, we can operate in two ways: The first way is to re-execute all the transactions in the blockchain in the EVM and trace all the calls. This process is extremely expensive [49] from a computational point of view. The alternative is to scan the Event log looking for events that emit a Transfer event. The second way is much faster and we estimate that it loses only 12% of the total number of tokens created by internal transactions. Moreover, the missing tokens are tokens that have never been used, traded, or transferred and are thus of little importance for our study (we discuss in detail the impact of this choice in Section 10). So, we parse all the logs of both blockchains, searching for smart contracts that emit a Transfer event compliant with the ERC-20 (resp. BEP-20) interface. Then, we use EtherScan [31] and BscScan [32] to retrieve the transactions that created these smart contracts and all the needed information.
At the end of these two steps, we have a collection of 3,087,274 and 4,534,599 smart contracts extracted from BSC and Ethereum, respectively. For each of them, we store the following information: The address of the contract, the block number in which the smart contract has been generated, the block in which the smart contract emits the last event, the EOA that deployed the smart contract or in the case of internal transactions the EOA address that triggers the first smart contract, the amount of gas used, the cost of the gas unit (gas price), the bytecode of the smart contract, and if the smart contract has been deployed by an EOA or through an internal transaction.
@@ 4.1.2 Token Identification
Smart contracts are not only used to create tokens, and not all smart contracts that emit a Transfer event are tokens (e.g., NFT contracts). Thus, we need to identify which of the retrieved smart contracts are compliant with ERC-20 (resp. BEP-20). Unfortunately, this is not a trivial task, and in the last years several works [21, 53, 15, 14, 27], attempted to face this problem with several approaches that we describe in Section 9. For our analysis, we follow the approach proposed by  [53, 15] that leverage the bytecode of smart contracts.
According to the Solidity specification [35], in the bytecode, smart contracts’ methods are identified by signatures that consist of the first 4 bytes of the Kekkack-256 hash of the method name and parameters’ type. Thus, to verify if a bytecode of a retrieved smart contract represents an ERC-20 (resp. BEP-20) compliant token, we verify if it contains at least all the signatures of the ERC-20 (resp. BEP-20) mandatory methods. Tab. 5 in the Appendix shows the signature of the mandatory and optional methods of the ERC-20 and BEP-20 interfaces.
Of the 4,534,599 smart contracts’ bytecode retrieved on the Ethereum blockchain, we find that 389,348 (8.5%) are ERC-20 tokens compliant, and 381,551 (98%) of them also implement the optional functions of the ERC-20 interface. Instead, in the BNB Smart Chain, we find that 1,887,484 out of 3,087,274 (61%) are BEP-20 compliant and, almost all of them also implement the optional methods of the BEP-20 interface. Although we found more smart contracts in Ethereum than in BSC (4,534,599 vs 3,087,274), there are many more compliant tokens in BSC (1,887,484) than in Ethereum (389,348). This discrepancy suggests that BSC may be a more interesting environment to study tokens and, possibly, their misuse.
Lastly, we retrieve all the information about the identified tokens, such as the name, the symbol, the number of decimals, and the total supply. We use the Ethereum-etl library and the Contract Application Binary Interface (ABI) [23]. The ABI is an interface between two program modules. It contains the specification for encoding/decoding methods and structures to interact with the machine code and interpret the results. Through the library, it is possible to instantiate smart contracts in an object-oriented manner and call its methods using an appropriate ABI. We instantiate the token contracts using an ABI that contains the specifications of ERC-20 (resp. BEP-20) methods and call the name(), symbol(), decimals(), totalSupply() methods.
At the end of the process, we have a dataset of ERC-20 (resp. BEP-20) tokens containing all the information about the smart contracts described in Section 4.1.1 and the related tokens. Table 1 shows the number of smart contracts on both blockchains.
@@ 4.2 Liquidity Pools Dataset
To create the Liquidity Pool dataset, we consider Uniswap, its forks, and other protocols that leverage its smart contracts.
@@
Uniswap has three main smart contracts: Factory, Pair and the Router. The Factory contract is responsible for creating the smart contract that handles the liquidity pool and the LP-tokens. Note that since the same smart contract handles the LP-token and the liquidity pool they have the same contract address. The Pair contract keeps track of the balances of the tokens in the pool and implements the AMM logic explained in Sec. 2. The Router contract offers the entry point to interact with the liquidity pools. Thus, interacting with the Router, it is possible to add or remove cryptocurrencies from a liquidity pool and swap tokens. Each of these contracts implements a set of Events that notify their status changes.
@@
To build our datasets, we parse the Event log of the Ethereum and BSC blockchains. Following there are the events we look for and a brief description:
@@
  • PairCreated: This event is fired by the Factory contract each time a new liquidity pool is created. We find 972,127 and 70,581 PairCreated events emitted in BSC and Ethereum, respectively. From the event, we can obtain the transaction hash, the block of the creation of the liquidity pool, the address that created the liquidity pool, the address of the liquidity pool, and the addresses of the two tokens (the pair of the liquidity pool), the gas used and the price paid per gas.
    Analyzing the address that fired the event and looking online for notable smart contract addresses, it is possible to have a rough idea of the diffusion of the Uniswap forks in the blockchains. In BSC, we find that PancakeSwap created most liquidity pools, with 941,220 emitted events (96.8%), followed by ApeSwap [4] (3,265 events), BakerySwap [6] (2,418 events) and Mdex [38] (1,602 events). In Ethereum, Uniswap emitted 65,098 events (92.2%), while the SushiSwap [12] Factory contract, a popular alternative to Uniswap on Ethereum, 2,637 (3%).
@@
  • Mint & Burn: The Pair contract emits a Mint (or Burn) Event each time a LP-token is minted (or burned). This occurs every time a liquidity provider adds (or removes) tokens into a liquidity pool. Analyzing these events, we can obtain the transaction hash and the block of the Mint (Burn) Event, the address of the liquidity pool, the address that added (removed) the liquidity, the number of LP-tokens minted (burned), the gas used, and price paid for the gas. We find 26,972,454 Mint events and 9,820,309 Burn events in BSC, and 1,911,918 Mint events and 1,067,841 Burn events in Ethereum.
@@
  • Swap: This event is fired by the Pair contract each time a user swaps tokens in a liquidity pool. From the event, we obtain all the information related to the swap as: The transaction hash, the block in which the swap occurs, the address of the liquidity pool used, the address that performs the swap, the number of tokens swapped, the gas used and the gas price. We find 750,508,160 events in BSC and 82,447,051 events in Ethereum.
@@
Moreover, we complete our dataset by collecting for each smart contract the block number in which it emits the last event. Tab. 2 describes the final dataset.
@@
LP-tokens are ERC-20 (resp. BEP-20) compliant tokens, and they are already present in our Tokens Dataset. However, our goal is to study standard token and liquidity pools separately. Thus, as the final step, we get rid of the information related to the LP-tokens from the Tokens Dataset. The last line on Tab. 1 reports the number of tokens after getting rid of the LP-tokens.
@@ 5 The Lifetime of Tokens
Our data collection revealed a surprisingly high number of tokens and liquidity pools in Ethereum and BSC. Services like CoinGecko [16] or CoinmarketCap [17] list about 13,000 cryptocurrencies on 602 centralized and decentralized exchanges in total. So, it is unclear what is the role of the large majority of tokens in the blockchain ecosystem.
Figure 1: Lifetime of tokens and liquidity pools in BSC and Ethereum.
@@
Figure 1 shows the CDF of tokens’ lifetime and liquidity pools’ lifetime in Ethereum (blue lines) and BSC (yellow lines). Tokens and liquidity pools are shown with solid and dashed lines, respectively. The slope of the curves tells that the lifetime of the tokens in BSC are generally shorter than the lifetime of the tokens in Ethereum. Consider that BSC is a young blockchain, with slightly more than two years of activity (released on 2020-04-20), while Ethereum is more than seven years old (released on 2015-07-30). The longevity of Ethereum is also visible by the long tail of its tokens in the CDF. Nonetheless, it is apparent that Ethereum hosts tokens that tend to be more solid and long-lasting. This difference is smaller when we look at liquidity pools. Indeed, PancakeSwap, that handles about 97% of the liquidity pools in BSC, was born only four months after the release of Uniswap V2. The reader can also observe that, in general, the lifetime of the liquidity pools is shorter than the lifetime of the tokens. This is not surprising. Indeed, a liquidity pool needs both tokens involved to be created in advance. Thus, the lifetime of a liquidity pool can be maximum as long as the lifetime of the younger token.
From the CDF, we can also note a few additional interesting facts, particularly when looking at the first 24 hours of the life of tokens and liquidity pools.
@@
A significant fraction of tokens are never active. Looking at the zoomed image in the center of Figure 1 (b), it is possible to see that a significant fraction of tokens have a lifetime of zero, meaning that the token is active only in 1 block when it was created. This phenomenon is more common in Ethereum, with 104,836 out of 323,863 (32.4%) tokens that belong to this category, against 167,318 out of 1,078,016 (15.5%) in BSC. In the following, we refer to the tokens that last only one block as 1-block tokens, while to the other tokens as active tokens. We find 910,698 and 219,027 active tokens in BSC and Ethereum, respectively. Table 3 succinctly reports on these statistics.
@@
A large part of active tokens have an extremely short lifetime. Figure 1 (b) shows that about 60% of the tokens in BSC and Ethereum have a lifetime shorter than one day. We refer to these tokens as 1-day tokens. If we consider only active tokens, we find that 471,385 (51.7%) of all the active BSC tokens and 82,542 (37.7%) of all the Ethereum active are 1-day tokens. Looking at the data at a higher granularity (Figure 1 (b)), we can note that the death ratio of BSC tokens is surprisingly high. BSC has approximately half of the 1-block tokens of Ethereum, about the same proportion of dead tokens after 60 minutes, and a significantly larger proportion of dead tokens after the first 4 hours. As we can see in Figure 1 (c), the first four hours of token life are crucial also in Ethereum.
@@
Almost all the BSC tokens with short lifetime have a liquidity pool. Here, we find one of the main differences between BSC and Ethereum. 414,936 out of 471,385 (88%) active tokens with a lifetime shorter than one day in BSC have a liquidity pool. In Ethereum, only 33% (26,817). It seems that in BSC, the liquidity pool is the main reason to create a token.
Table 3: Summary of 1-day and 1-block tokens in the BNB Smart Chain and Ethereum.
Lifetime BNB Smart Chain Ethereum
@@
6 Who Creates Tokens?
In this section, we change perspective and explore who creates tokens. Retrieving the list of creator addresses from our token dataset, we find 144,795 and 464,095 different addresses that create at least one token, respectively in Ethereum and BSC. Comparing these numbers with the total number of cumulative unique addresses in Ethereum (189,858,744) and BSC (140,522,222) [^2] Data retrieved from Etherscan and BSCscan respectively, we see that they represent only a very small fraction of the addresses, the 0.07% in Ethereum and 0.33% in BSC. Figure 2 shows the distribution of the number of tokens created by addresses in Ethereum and BSC. The first thing to notice is that the two distributions are extremely similar. The large majority of these addresses (70%) create only one token, as we can see in the zoomed image on the bottom right corner of Figure 2. 95% of addresses create 5 tokens or less and just 1% of addresses create more than 18 tokens.
Figure 2: Distribution of the number tokens created by the addresses that create at least one token in BSC and Ethereum. For the sake of visualization, the CDF is cut at 100 tokens. The maximum number of tokens created are 17,936 in BSC and 1,740 in Ethereum.
However, we can gather further insights by plotting the same data differently.
A small fraction of addresses creates a disproportionate amount of tokens. Figure 3 shows the CDF of tokens created by fraction of addresses. From the figure, we can see that although 70% of addresses create just 1 token, the total amount of tokens created by these addresses account for only 30% of the tokens on the two blockchains. And, more interestingly, we find that just 1% of the addresses creates 24.3% (262,023) of the tokens in BSC, and similarly, 1% of the addresses in Ethereum create 20.1% (67,838) of the tokens. These addresses create an average of 51 and 61 tokens in Ethereum and BSC, respectively. We will refer to these addresses as token spammers.
6.1 The Token Spammers

### v3

@@ 4.1.1 Gathering smart contracts
As a first step to building the Token dataset, we collect all the contract creation transactions issued by EOAs. As mentioned in Sec. 2, EOAs can deploy a smart contract by sending a contract creation transaction to the zero address. We process all the transactions in the considered time frame in BNB Smart Chain (2.6 billion transactions) and in Ethereum (1.4 billion transactions). We collect 2,195,399 and 4,420,389 contract creation transactions respectively.
However, tokens can also be created by a smart contract itself. Indeed, it could be the case that an EOA calls a smart contract method, and its execution generates a new ERC-20 (or BEP-20) compliant smart contract. In this case, the token is created with a so-called internal transaction. Despite the name, internal transactions are not real transactions, but rather calls performed by smart contracts. These kinds of transactions are stored off-chain—they are not visible simply parsing the blockchain.
To track the tokens created by internal transactions, we can operate in two ways: The first way is to re-execute all the transactions in the blockchain in the EVM and trace all the calls. This process is extremely expensive [57] from a computational point of view. The alternative is to scan the Event log looking for events that emit a Transfer event. The second way is much faster and we estimate that it loses only 12% of the total number of tokens created by internal transactions. Moreover, the missing tokens are never been used, traded or transferred, and are thus of little importance for our study (we discuss in detail the impact of this choice in Sec. 11). So, we parse all the logs of both blockchains, searching for smart contracts that emit a Transfer event compliant with the ERC-20 (resp. BEP-20) interface. Then, we use Etherscan [36] and BscScan [37] to retrieve the transactions that created these smart contracts and all the information.
At the end of these two steps, we have a collection of 3,087,274 and 4,534,599 smart contracts extracted from BSC and Ethereum, respectively. For each of them, we store the following information: The address of the contract, the block number in which the smart contract has been generated, the block in which the smart contract emits its last event, the EOA that deployed the smart contract or in the case of internal transactions the EOA address that triggers the first smart contract, the amount of gas used, the cost of the gas unit (gas price), the bytecode of the smart contract, and if the smart contract has been deployed by an EOA or through an internal transaction.
@@ 4.1.2 Token identification
Smart contracts are not only used to create tokens, as well as not all smart contracts that emit a Transfer event are tokens (e.g., NFT contracts). Thus, we need to identify which of the retrieved smart contracts are ERC-20 (resp. BEP-20) compliant. Unfortunately, this is not a trivial task, and in the last years several works [22, 63, 14, 13, 28], attempted to face this problem with several approaches that we describe in Sec. 10. For our analysis, we follow the approach proposed by  [63, 14] that leverage the bytecode of smart contracts.
According to the Solidity specification [40], in the bytecode, smart contract’s methods are identified by signatures that consist of the first 4 bytes of the Kekkack-256 hash of the method name and parameters’ type. Thus, to verify if a bytecode of a retrieved smart contract represents an ERC-20 (resp. BEP-20) compliant token, we verify if it contains at least all the signatures of the ERC-20 (resp. BEP-20) mandatory methods. Tab. 5 in the Appendix shows the signature of the mandatory and optional methods of the ERC-20 and BEP-20 interfaces.
Of the 4,534,599 smart contracts’ bytecodes retrieved on the Ethereum blockchain, we find that 389,348 (8.5%) are ERC-20 tokens compliant, and 381,551 (98%) of them also implement the optional functions of the ERC-20 interface. Instead, on the BNB Smart Chain, we find that 1,887,484 out of 3,087,274 (61%) are BEP-20 compliant, and, as for Ethereum, almost all of them also implement the optional methods of the BEP-20 interface. Although we found more smart contracts on Ethereum than in BSC (4,534,599 vs. 3,087,274), there are many more compliant tokens in BSC (1,887,484) than in Ethereum (389,348). This discrepancy suggests that BSC may be a more interesting environment to study tokens and, possibly, their misuse.
Lastly, we retrieve all the information about the identified tokens such as the name, the symbol, the number of decimals, and the total supply. To do so, we use the Ethereum-etl library and the Contract Application Binary Interface (ABI) [24]. The ABI is an interface between two program modules. It contains the specification for encoding/decoding methods and structures to interact with the machine code and interpret the results. Through the library, it is possible to instantiate smart contracts in an object-oriented manner and call its methods using an appropriate ABI. We instantiate the token contracts using an ABI that contains the specifications of ERC-20 (resp. BEP-20) methods and call the name(), symbol(), decimals(), totalSupply() methods.
At the end of the process, we have a dataset of ERC-20 (resp. BEP-20) tokens containing all the information about the smart contracts described in Sec. 4.1.1 and the related tokens. Table 1 shows the number of smart contracts on both blockchains.
@@ 4.2 Liquidity Pools dataset
To create the Liquidity Pool dataset, we consider Uniswap, its forks, and the other protocols that leverage its smart contracts.
@@
Uniswap has three main smart contracts: Factory, Pair, and the Router. The Factory contract is responsible for creating the smart contract that handles the liquidity pool and the LP-tokens. The Pair contract keeps track of the balances of the tokens in the pool and implements the AMM logic explained in Sec. 2. The Router contract offers the entry point to interact with the liquidity pools. Thus, it is possible to swap tokens and add or remove cryptocurrencies from a liquidity pool by interacting with the Router. Each of these contracts implements a set of Events that notify their status changes.
@@
To build our datasets, we parse the Event log of the Ethereum and BSC blockchains. Following, we report the events we look for and a brief description:
@@
  • PairCreated: This event is fired by the Factory contract each time a new liquidity pool is created. We find 972,127 and 70,581 PairCreated events emitted on BSC and Ethereum, respectively. From the event, we can obtain the transaction hash, the block of the creation of the liquidity pool, the address that created the liquidity pool, the address of the liquidity pool, and the addresses of the two tokens (the pair of the liquidity pool), the gas used and the price paid per gas. Analyzing the address that fired the event, we find that almost all the liquidity pools of BSC are created in PancakeSwap (96.8%), and almost all the liquidity pools of Ethereum are created in Uniswap (92.2%).
@@
  • Mint & Burn: The Pair contract emits a Mint (or Burn) Event each time an LP-token is minted (or burned). This occurs whenever a liquidity provider adds (or removes) tokens into a liquidity pool. Analyzing these events, we obtain the transaction hash and the block of the Mint (Burn) Event, the address of the liquidity pool, the address that added (removed) the liquidity, the number of LP-tokens minted (burned), the gas used, and the price paid for the gas. We find 26,972,454 Mint events and 9,820,309 Burn events on BSC, and 1,911,918 Mint events and 1,067,841 Burn events on Ethereum.
@@
  • Swap: This event is fired by the Pair contract each time a user swaps tokens in a liquidity pool. From the event, we obtain all the information related to the swap: The transaction hash, the block in which the swap occurs, the address of the liquidity pool used, the address that performs the swap, the number of tokens swapped, the gas used and the gas price. We find 750,508,160 events on BSC and 82,447,051 events on Ethereum.
@@
Moreover, we complete our dataset collecting for each smart contract the block number in which it emits the last event. Tab. 2 describes the final dataset.
@@
Given that LP-tokens are ERC-20 (resp. BEP-20) compliant tokens, they are already present in our Tokens Dataset. However, our goal is to study standard tokens and liquidity pools separately. Thus, as the final step, we get rid of the information related to the LP-tokens from the Tokens Dataset. The last line on Tab. 1 reports the number of tokens after we get rid of the LP-tokens.
@@ 5 The Lifetime of tokens
Our data collection revealed a surprisingly high number of tokens and liquidity pools on Ethereum and BSC. Services like CoinGecko [16] or CoinmarketCap [18] list about 13,000 cryptocurrencies on 602 centralized and decentralized exchanges. Therefore, it is unclear what is the role of the large majority of tokens in the blockchain ecosystem.
Figure 1: Lifetime of tokens and liquidity pools on BSC and Ethereum.
@@
Fig. 1 shows the CDF of tokens’ lifetime and liquidity pools’ lifetime on Ethereum (blue lines) and BSC (yellow lines). Tokens and liquidity pools are shown with solid and dashed lines, respectively. The slope of the curves tells that the lifetime of the tokens in BSC is generally shorter than the lifetime of the tokens in Ethereum. Consider that BSC is a young blockchain, with slightly more than two years of activity (released on 2020-04-20), while Ethereum is more than seven years old (released on 2015-07-30). The longevity of Ethereum is also visible by the long tail of its tokens in the CDF. Nonetheless, it seems that Ethereum’s tokens that tend to be more solid and long-lasting. This difference is smaller when we look at liquidity pools. Indeed, PancakeSwap, which handles about 97% of the liquidity pools in BSC, was born only four months after the release of Uniswap V2. From the CDF, we can also note a few additional interesting facts, particularly when we look at the first 24 hours of the life of tokens and liquidity pools.
@@
A significant fraction of tokens is never active. Looking at the zoomed image in the center of Fig. 1 (b), it is possible to see that a significant fraction of tokens have a lifetime of length zero, meaning that the token is active only in one block, when it was created. This phenomenon is more common in Ethereum, with 104,836 out of 323,863 (32.4%) tokens that belong to this category, against 167,318 out of 1,078,016 (15.5%) in BSC. In the following, we refer to the tokens that last only one block as 1-block tokens, while to the other tokens as active tokens. We find 910,698 and 219,027 active tokens on BSC and Ethereum, respectively. Table 3 succinctly reports on these statistics.
@@
A large part of active tokens has an extremely short lifetime. Fig. 1 (c) shows that about 60% of the tokens in BSC and Ethereum have a lifetime shorter than one day. We refer to these tokens as 1-day tokens. Considering only active tokens, we find that 471,385 (51.7%) of all the active BSC tokens and 82,542 (37.7%) of all the Ethereum active tokens are 1-day tokens. Looking at the data at a higher granularity (Fig. 1 (b)), we can note that the death ratio of BSC tokens is surprisingly high. Proportionally, BSC has approximately half of the 1-block tokens of Ethereum, about the same proportion of dead tokens after 60 minutes, and a significantly larger proportion of dead tokens after the first 4 hours. As we can see in Fig. 1 (c), the first four hours of token life are also crucial in Ethereum.
@@
Almost all the BSC tokens with short lifetimes have a liquidity pool. Here, we find one of the main differences between BSC and Ethereum. 468,556 out of 471,385 (94.8%) active tokens with a lifetime shorter than one day in BSC have a liquidity pool. In Ethereum, only 33.1% (27,346). It seems that on BSC the liquidity pool is the main reason for creating a token.
Table 3: Summary of 1-day and 1-block tokens for BSC and Ethereum.
Lifetime BSC Ethereum
@@
6 Token spammers
In this section, we change perspective and explore who creates tokens. Retrieving the list of creator addresses from our token dataset, we find 144,795 and 464,095 different addresses that create at least one token, respectively, in Ethereum and BSC. Comparing these numbers with the total number of cumulative unique addresses on Ethereum (189,858,744) and BSC (140,522,222) [^3] Data retrieved from Etherscan and BSCscan respectively, we see that they represent only a very small fraction of the addresses, the 0.07% in Ethereum and 0.33% in BSC. Fig. 2 shows the distribution of the number of tokens created by addresses in Ethereum and BSC. The first thing to notice is that the two distributions are extremely similar. The large majority of these addresses (70%) create only one token, as we can see in the zoomed image on the bottom right corner of Fig. 2. 95% of addresses create five tokens or less, and just 1% of addresses create more than 18 tokens.
Figure 2: Distribution of the number of tokens created by the addresses that create at least one token in BSC and Ethereum. For the sake of visualization, the CDF is cut at 100 tokens.
A small fraction of addresses creates a disproportionate amount of tokens. Fig. 3 shows the CDF of tokens created by fraction of addresses. From the figure, we can see that although 70% of addresses create just one token, the total amount of tokens created by these addresses account for only 30% of the tokens on the two blockchains. And more interestingly, we find that just 1% of the addresses creates 24.3% (262,023) of the tokens in BSC, and similarly, 1% of the addresses in Ethereum create 20.1% (67,869) of the tokens. These addresses create an average of 51 and 61 tokens in Ethereum and BSC, respectively. We will refer to these addresses as token spammers.